# Sampling 3D Molecular Conformers with Diffusion Transformers

### v1

@@ A.2 Compute and Training Times
All our models on GEOM QM9 are trained for 250 epochs, which requires 2 days of training on Nvidia H100 GPU for aPE and rPE models and almost 4 days for PE(3). Due to compute constraints, our non-equivariant (DiTMC+aPE) models are trained for 50 epochs on GEOM drugs. The equivariant models are trained for 10 epochs on GEOM drugs resulting in a similar compute budget for training in terms of GPU hours. Our maximal compute budget per model is 9 days on a single Nvidia H100 GPU.
@@
Table A5: Architectural details as well as training and inference time (in ms) for different PE strategies on QM9. Times are measured with batch size 128 on a single Nvidia H100 GPU. $T$ means number of transformer layers, $n_{\text{heads}}$ number of heads, $d_{\text{head}}$ number of features per head in the attention update, and $T_{MGN}$ number of layers for the conditioning mesh graph net. Thus, total feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$.
@@
Appendix B Architecture Details
Tab. A5 shows the architectural details for our base and large model variants. Tab. A6 contains details about the MLPs we use throughout our architecture.
Table A6: Architecture details for MLPs used in the model. The feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$ where $n_{\text{head}}$ is the number of heads and $d_{\text{head}}$ is the number of features per head. The use of gated GELU ensures equivariance (see section Sec. B.4).
Name Layers Hidden Dim Out Dim Activation Input
DiT Block 2 $4H$ $H$ GELU Tokens
SO(3) DiT Block 2 $4H$ $H$ gated GELU Tokens
Bond pair 2 $H$ $H$ SiLU $\boldsymbol{c}_{ij}^{\mathcal{G}}$
Time embedding 2 $H$ $H$ SiLU $t\in[0,1]$
Shortest-hop embedding 2 $H$ $H$ SiLU Hop distance
aPE embedding 2 $H$ $H$ SiLU Abs. ositions $\vec{r}_{i}$
rPE embedding 2 $H$ $H$ SiLU Rel. positions $\vec{r}_{ij}$
GNN embedding 2 $H$ $H$ SiLU Node/edge features
We first describe the building blocks of our non-equivariant DiTMC implementation and then its equivariant counterpart. Both architectures rely on conditioning tokens $\mathcal{C}$, for the time $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, per-node $\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}$, per bond-pair $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ and for all-pairs $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$. See main text section 4.1 for more details. Following Ref. [13], we use adaptive layer norm (AdaLN) and adaptive scale (AdaScale) to include per-node conditioning based on time $t$ and molecular graph information. To that end, we construct conditioning tokens
$\boldsymbol{c}_{i}=\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}.$ (A28)
B.1 Non Equivariant DiT
In the non-equivariant DiTMC formulations with aPE and rPE, we have the following set of tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$.
B.1.1 Adaptive Layer Normalization and Adaptive Scale
@@
where $\boldsymbol{W}\in\mathbb{R}^{6H\times H}$ and the output is split into six equally sized vectors $\boldsymbol{\alpha}_{1i},\boldsymbol{\beta}_{1i},\boldsymbol{\gamma}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\beta}_{2i},\boldsymbol{\gamma}_{2i}\in\mathbb{R}^{H}$. The weight matrix $\boldsymbol{W}$ is initialized to all zeros, such that “AdaLN” behaves like identity at initialization. “AdaScale” damps all input tokens to zero at initialization such that the whole DiT block behaves like the identity function at initialization.
B.1.2 Readout
@@
B.2 SO(3) Equivariant DiT
Following the notation in Ref. [60], we denote $\mathrm{SO}(3)$-equivariant tokens as =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\mathbb{R}^{(L+1)^{2}\times H}\}$, where $L$ denotes the maximal degree of the spherical harmonics. We denote the features corresponding to the $\ell$-th degree as $\boldsymbol{h}_{i}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$, where the $(2\ell+1)$ entries corresponds to the orders $m=-\ell,\dots,+\ell$ per degree $\ell$. We refer the reader to Ref. [60] for an in-depth introduction into equivariant features. For all our experiments we use maximal degree of $L=1$.
B.2.1 Adaptive Layer Normalization and Adaptive Scale
In the $\mathrm{SO}(3)$-equivariant case, we define adapted version of AdaLN and AdaScale, preserving equivariance. Our equivariant formulation of AdaLN is given as
@@ B.2.1 Adaptive Layer Normalization and Adaptive Scale
where EquivLN is the equivariant formulation of layer normalization following Ref. [36] without trainable per-degree scales and LN is standard layer normalization without trainable scale and bias. Scaling vectors $\boldsymbol{\alpha}^{(\ell)}\in\mathbb{R}^{1\times H}$ are defined per degree $\ell$, such that input scaling vectors are tensors $\boldsymbol{\alpha}\in\mathbb{R}^{(L+1)\times H}$. Bias vectors are only defined for the invariant ($\ell=0$) component of the tokens, since adding a non-zero bias to components with $\ell>0$ would lead to a non-equivariant operation (the bias does not transform under rotations). The element wise multiplication between $(1+\boldsymbol{\alpha}^{(\ell)})\in\mathbb{R}^{1\times H}$ and tokens $h^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$ is “broadcasted” along the degree-axis. For $L=0$, Eq. A33 reduces to the standard adaptive layer normalization.
@@
B.2.2 Equivariant positional embeddings
@@
$\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\boldsymbol{R}\vec{r}_{ij})$ $=\boldsymbol{\phi}_{\ell c}(||\boldsymbol{R}\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\boldsymbol{R}\hat{r}_{ij})$ (A39)
$=\boldsymbol{\phi}_{\ell c}(||\vec{r}_{ij}||)\,\odot\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{Y}_{\ell}(\hat{r}_{ij}),$ (A40)
@@
B.3 Invariance of the Dot-Product
@@
B.3.1 Readout
B.4 Equivariant MLP
Standard MLPs do not preserve the equivariance of the tokens. However, it is possible to define an equivariant formulation for dense layers and so-called gated non-linearities which preserve equivariance. We use them to build equivariant MLPs for the node-wise refinement after the self-attention calculation. See e.g. Ref. [60] for more details.
Appendix C Sampling
To ensure $\mathrm{SE}(3)$ invariance of the probability path from an (approximately) $\mathrm{SO}(3)$-equivariant velocity predictor, we center the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$ as well as the prediction of DiTMC in each ODE step.
Algorithm 2 ODE Sampling
1:Model $\boldsymbol{x}_{1}^{\theta}$, Graph $\mathcal{G}$, steps $N>0$
2:
3:$t_{n}\leftarrow n/N$ for $n\in\{0,\ldots,N\}$
4:$\boldsymbol{x}_{0}\sim p_{\text{prior}}(\boldsymbol{x})$triangleright Sample prior.
5:$\boldsymbol{x}_{0}\leftarrow\mathrm{Center}(\boldsymbol{x}_{0})$
6:for $n\leftarrow 0$ to $N-1$ do
7:     $\Delta t\leftarrow t_{n+1}-t_{n}$triangleright Compute step size.
11:     $\boldsymbol{x}_{t_{n+1}}\leftarrow\boldsymbol{x}_{t_{n}}+\Delta t\cdot\boldsymbol{v}$triangleright Euler step.
12:end for
13:
14:return $\boldsymbol{x}_{1}$
Appendix D Implementation details
D.1 Data Preprocessing
For both GEOM-QM9 and GEOM-DRUGS, we use the first 30 conformers for each molecule with the lowest energies, i.e., highest Boltzmann weights. We use the train/test/val split from Geomol [65], using the same 1000 molecules for testing.
D.2 Input Featurization
Tab. A7 defines the features we use for each atom or bond. Each feature is one-hot encoded before being passed to the network.

### v2

@@ Appendix B Training
All models on GEOM-QM9 are trained for 250 epochs, which requires 2 days of training on Nvidia H100 GPU for aPE and rPE models and almost 4 days for PE(3). For GEOM-DRUGS, we fix the total compute budget per model to nine days on a single NVIDIA H100 GPU, due to computational constraints. Within this budget, we can train aPE-L and rPE-L for 50 epochs and PE(3)-L for 10 epochs. To ensure consistency within each PE strategy, the base (“B”) model are trained for the same number of epochs as the corresponding large (“L”) model.
Appendix C Sampling
To ensure $\mathrm{SE}(3)$ invariance of the probability path from an (approximately) $\mathrm{SO}(3)$-equivariant velocity predictor, we center the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$ and the prediction of DiTMC in each ODE step.
Algorithm 2 ODE Sampling
1:Model $\boldsymbol{x}_{1}^{\theta}$, Graph $\mathcal{G}$, steps $N>0$
2:
[…]
6:for $n\leftarrow 0$ to N1N-1 do
7:  $\Delta t\leftarrow t_{n+1}-t_{n}$ $\triangleright$ Compute step size.
11:  $\boldsymbol{x}_{t_{n+1}}\leftarrow\boldsymbol{x}_{t_{n}}+\Delta t\cdot\boldsymbol{v}$ $\triangleright$ Euler step.
12:end for
13:
14:return $\boldsymbol{x}_{1}$
Appendix D Architectural Details
In this section, we describe the architecture of DiTMC and the conditioning GNN in more detail.
All DiTMC models rely on conditioning tokens $\mathcal{C}$, for the time $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, per-atom $\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}$, and per atom-pair $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$. See main text Section 4.1 for more details. Following Ref. [peebles2023DiT], we use adaptive layer norm (AdaLN) and adaptive scale (AdaScale) to include per-atom conditioning tokens based on time $t$ and molecular graph information. To that end, we construct conditioning tokens
$\boldsymbol{c}_{i}=\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}.$ (A22)
Tab. A6 contains details on the MLPs used throughout our architecture, while Tab. A7 summarizes architectural details, as well as training and inference times for all DiTMC variants. Note that the number of attention heads in DiTMC+PE(3) is adjusted to match the total parameter count of the corresponding DiTMC+aPE and DiTMC+rPE variants, while all other hyperparameters are identical.
Table A6: Architectural details for MLPs used in the model. The feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$ where $n_{\text{head}}$ is the number of heads and $d_{\text{head}}$ is the number of features per head.
Name Layers Hidden Dim Out Dim Activation Input
DiTMC Block 2 $4H$ $H$ GELU Tokens
SO(3) DiTMC Block 2 $4H$ $H$
gated
GELU
Tokens
Bond pair 2 $H$ $H$ SiLU $\boldsymbol{c}_{ij}^{\mathcal{G}}$
Time embedding 2 $H$ $H$ SiLU $t\in[0,1]$
Shortest-hop embedding 2 $H$ $H$ SiLU Hop distance
aPE embedding 2 $H$ $H$ SiLU Abs. positions $\vec{r}_{i}$
rPE embedding 2 $H$ $H$ SiLU Rel. positions $\vec{r}_{ij}$
GNN embedding 2 $H$ $H$ SiLU Node/edge features
@@ Appendix D Architectural Details
Table A7: Architectural details for different PE strategies on GEOM-QM9 and GEOM-DRUGS. Times are measured on GEOM-QM9 with batch size 128 on a single Nvidia H100 GPU. $T$ means number of transformer layers, $n_{\text{heads}}$ number of heads, $d_{\text{head}}$ number of features per head in the attention update, and $T_{MGN}$ number of layers for the conditioning mesh graph net. Thus, total feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$.
@@
D.1 Non-Equivariant DiTMC
In the non-equivariant DiTMC formulations based on aPE and rPE, we have the following set of tokens $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$. Initial token representations are obtained via $\boldsymbol{h}_{i}=\textbf{e}(z_{i})+\boldsymbol{p}_{i}^{\text{aPE}}$ for aPE and $\boldsymbol{h}_{i}=\boldsymbol{e}(z_{i})$ for rPE, where $\boldsymbol{e}(z_{i})\in\mathbb{R}^{H}$ denotes a learnable embedding based on the atomic number $z_{i}\in\mathbb{N}_{+}$ of atom $i$ [frank2024fastandstable].
D.1.1 Self-Attention Operation
For ease of notation, we only describe self-attention with a single head, but employ multi-head attention [vaswani2017attention] with $n_{\text{heads}}$ heads in our experiments. All self-attention blocks rely on query, key and value vectors, which are obtained from the input tokens $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\ |\ \boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ as
$\boldsymbol{q}=\boldsymbol{W}_{q}\boldsymbol{h}\,,\hskip 25.0pt\boldsymbol{k}=\boldsymbol{W}_{k}\boldsymbol{h}\,,\hskip 25.0pt\boldsymbol{v}=\boldsymbol{W}_{v}\boldsymbol{h}\,$, (A23)
where $\boldsymbol{W}_{q},\boldsymbol{W}_{k},\boldsymbol{W}_{v}\in\mathbb{R}^{H\times H}$ are trainable weight matrices. We define a slightly modified similarity kernel
$\text{sim}(\boldsymbol{q},\boldsymbol{k},\boldsymbol{u})=\exp\Bigg(\frac{\boldsymbol{q}^{\intercal}\cdot(\boldsymbol{k}\,\odot\boldsymbol{u})}{\sqrt{H}}\Bigg),$ (A24)
where $\boldsymbol{u}\in\mathbb{R}^{H}$ is used to inject additional information, e.g., conditioning signals and/or positional embeddings, and $\odot$’ denotes element-wise multiplication.
For absolute and relative PEs, we slightly modify standard self-attention to allow injecting pair-wise information into the values in addition to using our modified similarity kernel
$\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}_{\text{Pair}}^{\mathcal{G}})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\boldsymbol{v}_{j}\odot\boldsymbol{u}_{ij})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,$, (A25)
where queries, keys, and values are obtained with Eq. A23 and the injected pair-wise information $\boldsymbol{u}_{ij}$ depends on the positional embedding strategy with
[…]
Here $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ are pair-wise graph conditioning tokens (see Eq. 7) and $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$ and $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$ are the absolute and relative PEs described above (see Eqs. 8 and 9).
D.1.2 Adaptive Layer Normalization and Adaptive Scale
@@
where $\boldsymbol{W}\in\mathbb{R}^{6H\times H}$ and the output is split into six equally sized vectors $\boldsymbol{\alpha}_{1i},\boldsymbol{\beta}_{1i},\boldsymbol{\gamma}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\beta}_{2i},\boldsymbol{\gamma}_{2i}\in\mathbb{R}^{H}$. The weight matrix $\boldsymbol{W}$ is initialized to all zeros, such that “AdaLN” behaves like identity at initialization. “AdaScale” damps all input tokens to zero at initialization such that the whole DiTMC block behaves like the identity function at initialization.
D.1.3 Readout
@@
D.2 SO(3) Equivariant DiTMC
Following the notation in Ref. [unkee3x], we denote $\mathrm{SO}(3)$-equivariant tokens as $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}\}$, where $L$ denotes the maximal degree of the spherical harmonics. We denote the features corresponding to the $\ell$-th degree as $\boldsymbol{h}_{i}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$, where the $(2\ell+1)$ entries corresponds to the orders $m=-\ell,\dots,+\ell$ per degree $\ell$. We refer the reader to Ref. [unkee3x] for an in-depth introduction into equivariant features. For the initial token representation $\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}$ we set $\boldsymbol{h}_{i}^{(0)}=\boldsymbol{e}(z_{i})$, where $\boldsymbol{e}(z_{i})\in\mathbb{R}^{1\times H}$ denotes a learnable embedding based on the atomic number $z_{i}\in\mathbb{N}_{+}$ for atom $i$ [frank2024fastandstable]. All higher order features $\boldsymbol{h}_{i}^{(\ell)}$ with degree $\ell>0$ are initialized with zero. For all our experiments we use maximal degree of $L=1$.
D.2.1 Self-Attention Operation
Our equivariant version of self-attention uses the same transformations for queries, keys and values like the non-equivariant counterpart, as well as the modified similarity kernel (see Appendix subsubsection D.2.1). However, to preserve all Euclidean symmetries throughout the network, every token must transform equivariantly. One way to achieve this is by separating out the rotational degrees of freedom, encoding them with irreducible representations of the rotation group $\mathrm{SO(3)}$. This introduces a “degree-axis” of size $(L+1)^{2}$, which encodes angular components of increasing order. The maximum degree $L$ is chosen to ensure high fidelity at a reasonable computational cost. For example, setting $L=1$ restricts the representation to scalars and vectors, as used in models like PaiNN [schutt2021equivariant] or TorchMDNet [tholke2022equivariant]. An $\mathrm{SO(3)}$-equivariant formulation of self-attention is then given as
$\text{ATT}_{\text{SO(3)}}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\hat{\boldsymbol{u}}_{ij}\otimes\boldsymbol{v}_{j})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,$, (A31)
where equivariant queries, keys and values can be calculated similarly to Eq. A23 and $\otimes$’ denotes a Clebsch-Gordan (CG) tensor product contraction [unkee3x]. The dot-product in the similarity measure is taken along both feature and degree axes, such that the overall update preserves equivariance (see Appendix Appendix F for details). Tokens and scaling vectors are calculated as
$\hskip 15.0pt\boldsymbol{u}_{ij}=\boldsymbol{\phi}(r_{ij})\odot\boldsymbol{c}_{ij}^{\mathcal{G}}\,,\hskip 15.0pt\hat{\boldsymbol{u}}_{ij}=\boldsymbol{p}_{ij}^{\text{PE(3)}}\odot\boldsymbol{c}_{ij}^{\mathcal{G}}\,$, (A32)
where $\boldsymbol{\phi}(r_{ij})\in\mathbb{R}^{(L+1)^{2}\times H}$ is a radial filter, and the element-wise products with the pair-wise conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{1\times H}$ are broadcast along the degree axis. Importantly, the $2\ell+1$ subcomponents of the radial filter for degree $\ell$ are obtained by repeating per-degree filter functions $\boldsymbol{\phi}_{\ell}(r_{ij})\in\mathbb{R}^{1\times H}$ along the degree axis to preserve equivariance (see also Eq. 10).
Since, standard MLPs do not preserve SO(3)-equivariance, we use equivariant MLPs for the node-wise refinement after the self-attention calculation via an equivariant formulation for dense layers and so-called gated non-linearities (see e.g. Ref. [unkee3x] for more details).
D.2.2 Adaptive Layer Normalization and Adaptive Scale
In the $\mathrm{SO}(3)$-equivariant case, we define an adapted version of AdaLN and AdaScale, preserving equivariance. Our equivariant formulation of AdaLN is given as
@@ D.2.2 Adaptive Layer Normalization and Adaptive Scale
where EquivLN is the equivariant formulation of layer normalization following Ref. [liao2024equiformerv] without trainable per-degree scales and LN is standard layer normalization without trainable scale and bias. Scaling vectors $\boldsymbol{\alpha}^{(\ell)}\in\mathbb{R}^{1\times H}$ are defined per degree $\ell$, such that input scaling vectors are tensors $\boldsymbol{\alpha}\in\mathbb{R}^{(L+1)\times H}$. Bias vectors are only defined for the invariant ($\ell=0$) component of the tokens, since adding a non-zero bias to components with $\ell>0$ would lead to a non-equivariant operation (the bias does not transform under rotations). The element wise multiplication between $(1+\boldsymbol{\alpha}^{(\ell)})\in\mathbb{R}^{1\times H}$ and tokens $h^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$ is “broadcasted” along the degree-axis. For $L=0$, Eq. A33 reduces to the standard adaptive layer normalization.
@@
D.2.3 Readout
D.3 Conditioning GNN
Our conditioning GNN directly operates on the molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to obtain graph-based information for atom-wise conditioning. First, we processes the molecular graph $\mathcal{G}$ to derive node and edge input features, initially represented as one-hot vectors (a full list of features is provided in Tab. A8). These features then are projected into a shared latent space using two-layer multilayer perceptrons (MLPs). Finally, we employ a GNN architecture inspired by the processor described in the MeshGraphNet (MGN) framework [pfaff2020learning], which refines the features via message passing.
The conditioning GNN maintains and updates both node and edge representations across multiple layers. Each message passing block consists of two main steps: first, the edge representations are updated based on the current edge representations and the representations of the connected nodes:
The final output of the described conditioning GNN is a set of node embeddings per atom, and a set of edge embeddings per bond. We use the final node embeddings as atom-wise graph conditioning tokens in DiTMC as detailed in Sec. 4.1. We want to highlight that one can also use the final edge embedding as pair-wise graph conditioning tokens in DiTMC. We discuss this further in Appendix I.2.
Appendix E Equivariance Proof for Euclidean Positional Embeddings
@@
[…]
@@
Appendix F Invariance Proof for the Dot-Product
@@
Appendix G Implementation details
G.1 Data Preprocessing
For both GEOM-QM9 and GEOM-DRUGS, we use the first 30 conformers for each molecule with the lowest energies, i.e., highest Boltzmann weights. We use the train/test/val split from Geomol [ganea2021geomol], using the same 1000 molecules for testing.
G.2 Input Featurization
Tab. A8 defines the features we use for each atom or bond. Each feature is computed using RDKit [landrum2013rdkit] and one-hot encoded before being passed to the network.